A) How large a sample should be taken if the population mean is to be estimated with 99% confidence to within $82? The population has a standard deviation of $906. (Round your answer up to the next whole number.)

B) Assume that z is the test statistic. (Give your answers correct to two decimal places.)

(a) Calculate the value of z for Ho: μ = 51, σ = 4, n = 39, x = 48.6.


(b) Calculate the value of z for Ho: μ = 20, σ = 3.9, n = 77, x = 21.4.


(c) Calculate the value of z for Ho: μ = 138.5, σ = 3.8, n = 19, x = 141.02.


(d) Calculate the value of z for Ho: μ = 815, σ = 42.9, n = 65, x = 798.4.
Answer #1

Question-A: Given that E-82 and σ-906 The ZCritical value at the 99% confidence level is: 2.576 (From normal tables) Calculate the sample size 2.576x9061 82 809.959 - 810 Hence, the required sample size is 810.

Question-B a) Calculate the value of Z for Ho : μ 51,0 4; n 39 and x 48.6 48.6-51 4/39 -3.74699879903904 3./5 Round to 2 deci

2 21.4-20 = 3.14998721598691 -3.15 (Round to 2 decimal place) Hence, the required test statistic is: 3.15. c) Calculate the value of Z for Ho : μ = 138.5; σ= 3.8: n= 19 and x = 141.02 141.02-138.5 3.8/V19 = 2.89063824676909 -2.89 (Round to 2 decimal place) Hence, the required test statistic is: 2.89

c) Calculate the value of Z for Ho : μ = 61 5:0 = 42.9: n = 65 and x = 798.4 798.4-815 42.9 65 =-3.1 1966150633464 --3.12 (Round to 2 decimal place) Hence, the required test statistic is: -3.12